Grade 12th passMechanics

moment of inertia of a ring and a disc of same mass about their diameter are same the ratio of their radii will be

santhosh
6 Years agoGrade 12th pass

1 Answer

Arun
6 Years ago
For Ring along diameter,I=MR^2/2 ...
for Disc along diameter,I=MR^2/4...
Given that I of ring is equal to I of disc,
MR^2/2=MR^2/4..
R^2(of ring)/R^2(of disc)=1/2.
R(of ring)/R(of disc)=1/√2